# SynthesizRR: Generating Diverse Datasets with Retrieval Augmentation

### v1

@@ Appendix C Influence of retrieval corpus on domain shift
Table 12: Effect of corpus-swapping for SynthesizRR 32-shot No-Retr-ICL. We generate only $4$k rows for AG News to reduce costs.
@@
RQ: Can SynthesizRR work effectively with relatively small corpora? In our main results §section 5, we assumed the existence of a large corpus, with minimum size of 0.9M documents. As noted, this corpus need not be unlabelled examples for our task; we were able to successfully generate customer reviews and product questions for Humor, Categoryand Polarity tasks, while retrieving from a corpus of product information (title and description).
@@
A potential problem with SynthesizRR is that corpuses of such massive size might be few in number. Thus, we compare the performance of SynthesizRR on CMU Movie Summary Bamman et al. (2013) and Yelp Dataset corpus Yelp (2022), which are between one to three orders of magnitude smaller than corpora used in Table 6. In Table 10 and Table 9, we see that SynthesizRR can perform suitably even with relatively small corpora (42k movie plots and 150k businesses). From the previous RQs, this suggests that the relevance of the corpus to the task is more important than the size of the corpus for the performance of SynthesizRR.
Appendix D Ablation: number of in-context examples
@@
Figure 7: Left: DeBERTa-v3L test accuracy ()($\uparrow$)( ↑ ), center: entity entropy ()($\uparrow$)( ↑ ), right: Mauve ()($\uparrow$)( ↑ ) for SynthesizRR Retr-ICL. We vary the number of in-context examples from 0 to 8. Teacher LLMs LLaMa-2 Chat 13B and Claude Instant-v1 are compared on 6 tasks: AG News, Hyperpartisan, ToI Headlines, Category, Humor and Polarity. We do not report Category 8-shot due to model failures.
@@
The use of in-context examples in the Retr-ICL variant of SynthesizRR leads to significant improvements in intrinsic and distillation metrics, as we saw in §5. Here, we do a deeper analysis on whether continually increasing the number of in-context examples yields a positive benefit.
@@
Appendix E Task inversion prompts and label verbalizations
@@ E.1 Hyperpartisan
Prompt E.1: Hyperpartisan FewGen In-context example:
“ Write a single news article using {label} . The written article should be 2 to 3 paragraphs long.
@@
“ Write a single news article using {label} . The written article should be 2 to 3 paragraphs long.
News Article: ”
Prompt E.2: Hyperpartisan SynthesizRR Retr-ICL In-context example:
“ News Article: {icl[article_retr]}
@@
“ News Article: {article_retr[k]}
@@
Rewritten Article: ”
Prompt E.3: Hyperpartisan SynthesizRR No-Retr-ICL In-context example:
“ Rewritten Article: {icl[gold_text]}
@@
“ News Article: {article_retr[k]}
@@
Rewritten Article: ”
E.2 ToI Headlines
@@
Prompt E.4: ToI Headlines FewGen In-context example:
“ Write a headline for a news article about {label} . The headline should be a single sentence.
@@
“ Write a headline for a news article about {label} . The headline should be a single sentence.
Headline: ”
Prompt E.5: ToI Headlines SynthesizRR Retr-ICL In-context example:
“ News Article: {icl[article_retr]}
@@
“ News Article: {article_retr[k]}
@@
Headline: ”
Prompt E.6: ToI Headlines SynthesizRR No-Retr-ICL In-context example:
“ Headline: {icl[article_retr]}
@@
“ News Article: {article_retr[k]}
@@ E.3 AG News
Prompt E.7: AG News FewGen In-context example:
“ Write a summary for a news article about {label} . The summary should be one or two short sentences.
@@
“ Write a summary for a news article about {label} . The summary should be one or two short sentences.
Summary: ”
Prompt E.8: AG News SynthesizRR Retr-ICL In-context example:
“News Article: {icl[article_retr]}
@@
“ News Article: {article_retr[k]}
@@
Summary: ”
Prompt E.9: AG News SynthesizRR No-Retr-ICL In-context example:
“ Summary: {icl[gold_text]}
@@
“ News Article: {article_retr[k]}
@@
Summary: ”
E.4 Category
In the Category dataset, we determine the product category from a review written by a user. For task inversion in SynthesizRR we must retrieve a product and prompt the frozen LLM to generate a user review within the same product-category as the retrieval query. Thus, we include a style-attribute to allow minor typos in the generation and restrict to a few sentences using a length-attribute.
@@ E.4 Category
Prompt E.10: Category FewGen In-context example:
“ Write a product review about a product which is in the category of {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
@@
“ Write a product review about a product which is in the category of {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: ”
Prompt E.11: Category SynthesizRR Retr-ICL In-context example:
“ Product details: {icl[product_retr]}
@@
“ Product details: {product_retr[k]}
@@
Review: ”
Prompt E.12: Category SynthesizRR No-Retr-ICL In-context example:
“ Review: {icl[gold_text]}
@@ E.5 Humor
Prompt E.13: Humor FewGen In-context example:
“ Write a short {label} question about a product on Amazon. Only include the question.
@@
“ Write a short {label} question about a product on Amazon. Only include the question.
Product Question: ”
Prompt E.14: Humor SynthesizRR Retr-ICL In-context example:
“ Product details: {icl[product_retr]}
Write a short {label} question about the above product on Amazon. Only include the question.
@@
“ Product details: {product_retr[k]}
Write a short {label} question about the above product on Amazon. Only include the question.
Product Question: ”
Prompt E.15: Humor SynthesizRR No-Retr-ICL In-context example:
“ Product Question: {icl[gold_text]}
@@
“ Product details: {product_retr[k]}
Write a short {label} question about the above product on Amazon. Only include the question.
Product Question: ”
E.6 Polarity
Polarity is a review-sentiment classification task. In SynthesizRR, the difficulty is increased as we must generate a review from a product. For task inversion, we prompt the LLM to generate a review which can have either positive or negative sentiment and include details from the retrieved product. As with Category, we allow typos and restrict the length to a few sentences using a length-attribute in the prompt.
@@ E.6 Polarity
Prompt E.16: Polarity FewGen In-context example:
“ Write a review about a product on Amazon which discusses {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
@@
“ Write a review about a product on Amazon which discusses {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: ”
Prompt E.17: Polarity SynthesizRR Retr-ICL In-context example:
“ Product details: {icl[product_retr]}
Write a review about the above product on Amazon which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
@@
“ Product details: {product_retr[k]}
Write a review about the above product on Amazon which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: ”
Prompt E.18: Polarity SynthesizRR No-Retr-ICL In-context example:
“ Review: {icl[gold_text]}
@@
“ Product details: {product_retr[k]}
Write a review about the above product on Amazon which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: ”
Appendix F Example generations
Here we showcase examples from the best-performing SynthesizRR approach (3-shot No-Retr-ICL using LLaMa-2 Chat 13B) for each of our 6 tasks. For brevity, we do not show the ICL examples, only the retrieved article and generated text.
@@ Appendix F Example generations
Appendix G Data Preprocessing
G.1 Datasets
Other datasets are not pre-processed.
G.2 Corpora

### v3

@@ Appendix D Influence of corpus on domain shift
Table 11: Effect of corpus-swapping for SynthesizRR 32-shot Non-RetrICL. We generate only $4$k rows for AG News to reduce costs.
@@
RQ: Can SynthesizRR work effectively with relatively small corpora? In our main results §5, we assumed the existence of a large corpus, with minimum size of 0.9M documents. As noted, this corpus need not be unlabelled examples for our task; we were able to successfully generate customer reviews and product questions for Humor, Category and Polarity tasks, while retrieving from a corpus of product information (title and description).
@@
A potential problem with SynthesizRR is that corpuses of such massive size might be few in number. Thus, we compare the performance of SynthesizRR on CMU Movie Summary Bamman et al. (2013) which is between one to three orders of magnitude smaller than other corpora in Table 6. In Table 7, we see that SynthesizRR can perform suitably even with such relatively small corpora (42k movie plots). From the previous RQs, this suggests that the relevance of the corpus to the task is more important than the size of the corpus for the performance of SynthesizRR.
Appendix E Dense vs sparse retrieval in SynthesizRR
Retriever AG. Hyp. ToI Cat. Hum. Pol. Avg.
(Size) ($8$K) ($2$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold 91.0 93.2 82.5 81.5 93.1 95.3 89.43
LLaMa2 Zero shot
Contr. 83.5 69.8 74.5 68.9 82.5 84.7 77.32
BM25 83.2 74.2 70.7 57.6 78.5 85.4 74.93
ClaudeV1 Zero shot
Contr. 83.9 72.3 71.8 66.8 62.1 88.7 74.29
BM25 83.2 57.2 69.8 53.7 73.9 91.8 71.60
LLaMa2 3-shot RetrICL
Contr. 83.0 78.5 73.3 72.4 90.2 91.0 81.38
BM25 82.1 77.9 71.9 65.4 87.5 87.4 78.69
ClaudeV1 3-shot RetrICL
Contr. 83.7 72.3 72.8 65.4 83.4 91.3 78.16
BM25 83.0 73.5 70.0 52.4 82.4 90.7 75.34
Table 12: Test accuracy after distilling a DeBERTa-v3L student on a dataset generated by SynthesizRR. Retrieval is done using BM25 and Contriever. We use the same seed set and corpus as Table 2. We train 5 student models and report the mean accuracy, rerunning all 5 in case of std $\geq$ 6.0. The top two subsections consider zero-shot synthesis and bottom two considers 3-shot RetrICL variant. We bold the best result in each subsection. Contriever numbers are reproduced from Table 6.
So far, a single dense retriever (Contriever) has been used for the content sourcing step by using a bi-encoder approach (Lee et al., 2019; Chen et al., 2017). We embed both the input in-context covariate and each corpus document, and then rank results based on cosine similarity. In §5, we retrieved $k=500$ documents for each in-context example and after filtering, randomly sampled among these to produce a grounded set of documents on which we apply our task inversion strategy RetrICL.
In this section we explore how changing the retrieval model affects the content sourcing stage and its downstream effects. Keeping other parts of the process the same, we switch Contriever to BM25 Okapi (Robertson and Zaragoza, 2009), a popular sparse retrieval method. Dense retrievers like Contriever perform a semantic match between the query and document, whereas BM25 performs only a lexical match based on inverse term frequencies, with no understanding of semantics. Additionally, BM25 outputs a score which is an unbounded positive number, thus we are unable to use meaningful thresholds to bound the similarity in our RetrICL approach. Instead, we construct the RetrICL in-context set using the top-2 retrieved contexts for each ICL example and without applying the filter.
We expect that picking semantically similar information is more important to SynthesizRR since we include a task inversion step, which intends to change the tone and lexical structure of the text while preserving its semantics. Thus, we want contexts which are semantically related to Gold data, to which we can apply stylistic or formatting transformations using a task-inversion prompt to bring it closer to Gold.
Surprisingly, in Table 7 we see that while intrinsic diversity from BM25-retrieved documents is often worse than Contriever, they both generate equally human-like text. However, comparing the DeBERTa-v3L accuracy of Contriever and BM25in Table 12, we see that a strong student model trained on a dataset obtained from the dense-retrieved document set consistently outperforms the sparse retriever BM25, which might be due to the filtering step we introduce in RetrICL. This filtering step leads to a reduction in mislabelling stemming from retrieving contexts that belong do a different class. Due to this, we conclude that dense retrieval models are potentially more suitable for SynthesizRR.
Appendix F Varying number of in-context examples in RetrICL
@@
Figure 7: Left: DeBERTa-v3L test accuracy ()($\uparrow$)( ↑ ), center: entity entropy ()($\uparrow$)( ↑ ), right: Mauve ()($\uparrow$)( ↑ ) for SynthesizRR RetrICL. We vary the number of in-context examples from 0 to 8. Teacher LLMs LLaMa-2 Chat 13B and Claude Instant-v1 are compared on 6 tasks: AG News, Hyperpartisan, ToI Headlines, Category, Humor and Polarity. We do not report Category 8-shot due to model failures.
@@
The use of in-context examples in the RetrICL variant of SynthesizRR leads to significant improvements in intrinsic and distillation metrics, as we saw in §5. Here, we do a deeper analysis on whether continually increasing the number of in-context examples yields a positive benefit.
@@
Appendix G Task inversion prompts and label verbalizations
@@ G.1 Hyperpartisan
Prompt G.1: Hyperpartisan FewGen In-context example:
Write a single news article using {label} . The written article should be 2 to 3 paragraphs long.
@@
Write a single news article using {label} . The written article should be 2 to 3 paragraphs long.
News Article:
Prompt G.2: Hyperpartisan SynthesizRR RetrICL In-context example:
News Article: {icl[article_retr]}
@@
News Article: {article_retr[k]}
@@
Rewritten Article:
Prompt G.3: Hyperpartisan SynthesizRR Non-RetrICL In-context example:
Rewritten Article: {icl[gold_text]}
@@
News Article: {article_retr[k]}
@@
Rewritten Article:
G.2 ToI Headlines
@@
Prompt G.4: ToI Headlines FewGen In-context example:
Write a headline for a news article about {label} . The headline should be a single sentence.
@@
Write a headline for a news article about {label} . The headline should be a single sentence.
Headline:
Prompt G.5: ToI Headlines SynthesizRR RetrICL In-context example:
News Article: {icl[article_retr]}
@@
News Article: {article_retr[k]}
@@
Headline:
Prompt G.6: ToI Headlines SynthesizRR Non-RetrICL In-context example:
Headline: {icl[article_retr]}
@@
News Article: {article_retr[k]}
@@ G.3 AG News
Prompt G.7: AG News FewGen In-context example:
Write a summary for a news article about {label} . The summary should be one or two short sentences.
@@
Write a summary for a news article about {label} . The summary should be one or two short sentences.
Summary:
Prompt G.8: AG News SynthesizRR RetrICL In-context example:
News Article: {icl[article_retr]}
@@
News Article: {article_retr[k]}
@@
Summary:
Prompt G.9: AG News SynthesizRR Non-RetrICL In-context example:
Summary: {icl[gold_text]}
@@
News Article: {article_retr[k]}
@@
Summary:
G.4 Category
In the Category dataset, we determine the product category from a review written by a user about products on a major e-commerce website. For task inversion in SynthesizRR we must retrieve a product and prompt the frozen LLM to generate a user review within the same product-category as the retrieval query. Thus, we include a style-attribute to allow minor typos in the generation and restrict to a few sentences using a length-attribute.
@@ G.4 Category
Prompt G.10: Category FewGen In-context example:
Write a product review about a product which is in the category of {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
@@
Write a product review about a product which is in the category of {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.11: Category SynthesizRR RetrICL In-context example:
Product details: {icl[product_retr]}
@@
Product details: {product_retr[k]}
@@
Review:
Prompt G.12: Category SynthesizRR Non-RetrICL In-context example:
Review: {icl[gold_text]}
@@ G.5 Humor
Prompt G.13: Humor FewGen In-context example:
Write a short {label} question about a product. Only include the question.
@@
Write a short {label} question about a product. Only include the question.
Product Question:
Prompt G.14: Humor SynthesizRR RetrICL In-context example:
Product details: {icl[product_retr]}
Write a short {label} question about the above product. Only include the question.
@@
Product details: {product_retr[k]}
Write a short {label} question about the above product. Only include the question.
Product Question:
Prompt G.15: Humor SynthesizRR Non-RetrICL In-context example:
Product Question: {icl[gold_text]}
@@
Product details: {product_retr[k]}
Write a short {label} question about the above product. Only include the question.
Product Question:
G.6 Polarity
Polarity is a sentiment classification task for reviews of products on a major e-commerce website. In SynthesizRR, the difficulty is increased as we must generate a review from a product. For task inversion, we prompt the LLM to generate a review which can have either positive or negative sentiment and include details from the retrieved product. As with Category, we allow typos and restrict the length to a few sentences using a length-attribute in the prompt.
@@ G.6 Polarity
Prompt G.16: Polarity FewGen In-context example:
Write a review about a product which discusses {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
@@
Write a review about a product which discusses {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.17: Polarity SynthesizRR RetrICL In-context example:
Product details: {icl[product_retr]}
Write a review about the above product which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
@@
Product details: {product_retr[k]}
Write a review about the above product which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.18: Polarity SynthesizRR Non-RetrICL In-context example:
Review: {icl[gold_text]}
@@
Product details: {product_retr[k]}
Write a review about the above product which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
G.7 IMDb
IMDb is a review-sentiment classification task. As with other review tasks, in the task inversion step we prompt the LLM to generate a review in either positive or negative sentiment. The context used by SynthesizRR is the plotline of a movie from CMU Movie Summary. As with Category and Polarity, we allow typos and restrict the length to a few sentences using a length-attribute in the prompt.
Label Verbalization
positive what the reviewer liked about the movie
negative what the reviewer disliked about the movie
Table 19: Task inversion verbalizations for IMDb.
Prompt G.19: IMDb FewGen In-context example:
Write a review which discusses {label} . Include relevant details about the movie. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Write a review which discusses {label} . Include relevant details about the movie. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.20: IMDb SynthesizRR RetrICL In-context example:
Movie details: {icl[plotline_retr]}
Write a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Movie details: {plotline_retr[k]}
Write a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.21: IMDb SynthesizRR Non-RetrICL In-context example:
Review: {icl[gold_text]}
Prompt:
Movie details: {plotline_retr[k]}
Write a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
G.8 SST-2
SST-2 is another review-sentiment classification task, however the examples are partial sentences from movie reviews which were extracted such that they contain the sentiment-heavy phrases. This, during the task inversion we prompt the Teacher LLM to generate a partial review sentence in either positive or negative sentiment. The context used by SynthesizRR is the plotline of a movie from CMU Movie Summary. We allow typos and restrict the length to one sentence using a length-attribute in the prompt.
Label Verbalization
positive what the reviewer liked about the movie
negative what the reviewer disliked about the movie
Table 20: Task inversion verbalizations for SST-2.
Prompt G.22: SST-2 FewGen In-context example:
Write a single sentence from a review which discusses {label} . Include relevant details about the movie. The review should only be a single short sentence. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Write a single sentence from a review which discusses {label} . Include relevant details about the movie. The review should only be a single short sentence. Add very minor typos.
Review:
Prompt G.23: SST-2 SynthesizRR RetrICL In-context example:
Movie details: {icl[plotline_retr]}
Write a single sentence from a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Movie details: {plotline_retr[k]}
Write a single sentence from a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence. Add very minor typos.
Review:
Prompt G.24: SST-2 SynthesizRR Non-RetrICL In-context example:
Review: {icl[gold_text]}
Prompt:
Movie details: {plotline_retr[k]}
Write a single sentence from a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence. Add very minor typos.
Review:
Appendix H Example generations
Here we showcase examples from the best-performing SynthesizRR approach (3-shot Non-RetrICL using LLaMa-2 Chat 13B) for each of our 6 tasks. For brevity, we do not show the ICL examples, only the retrieved article and generated text.
@@ Appendix H Example generations
Appendix I Data Preprocessing
Country of headquarters No. articles Domain
Antigua 2.6K antiguaobserver.com
Azerbaijan 70.7K trend.az
Bangladesh 28.2K thedailystar.net
Barbados 5.1K caribbean360.com
Brazil 930 thebraziltimes.com
China 10.7K chinadigitaltimes.net, china.org.cn
Colombia 22.9K colombiareports.com, insightcrime.org
Costa Rica 18.9K ticotimes.net
Cuba 1.6K escambray.cu
Cyprus 13.2K cyprus-mail.com, dailyforex.com
Czech Republic 1.2K praguepost.com
Egypt 43 thedailynewsegypt.com
Estonia 21.2K err.ee
Ghana 5.2K ghanabusinessnews.com, modernghana.com
Guyana 70.2K stabroeknews.com
Hong Kong 5.6K asiasentinel.com, actionforex.com, hku.hk
India 886.5K mid-day.com, financialexpress.com, livemint.com, hindustantimes.com, indianexpress.com, mangalorean.com, vccircle.com, deccanchronicle.com, afaqs.com, bollywoodhungama.com, medianewsline.com, orissadiary.com, morungexpress.com, countercurrents.org, businessworld.in, governancenow.com, koimoi.com, milligazette.com, dayafterindia.com, truthdive.com, newstodaynet.com, centralchronicle.com, dalje.com, rtn.asia, realbollywood.com, mutiny.in
Indonesia 2K thejakartaglobe.com
Iran 7.2K tehrantimes.com
Israel 60.4K jewishpress.com, ynetnews.com, palestinechronicle.com, 972mag.com, defense-update.com
Jamaica 96.6K jamaica-gleaner.com
Japan 2.1K japantoday.com
Kenya 158.8K capitalfm.co.ke, nation.co.ke, theeastafrican.co.ke, standardmedia.co.ke, kbc.co.ke, businessdailyafrica.com
Kuwait 16.2K arabtimesonline.com, kuwaittimes.net
Lebanon 4.9K yalibnan.com
Macau 3.4K macaudailytimes.com.mo
Malawi 2.8K maravipost.com
Malaysia 30.5K malaysiakini.com, freemalaysiatoday.com, theborneopost.com
Misc. Africa 51 african-bulletin.com
Misc. Asia 30.9K eurasiareview.com
Namibia 20.2K newera.com.na
Nepal 2.2K thehimalayantimes.com
Nigeria 336.5K thenationonlineng.net, vanguardngr.com, thisdaylive.com, codewit.com, sunnewsonline.com, businessdayonline.com, pmnewsnigeria.com
Pakistan 274.1K nation.com.pk, dawn.com, tribune.com.pk, pakobserver.net, app.com.pk, dailytimes.com.pk, thefrontierpost.com, pakistankakhudahafiz.com, thenews.com.pk, pak1stanfirst.com, pakwatan.com
Palestine 655 intifada-palestine.com, paltelegraph.com
Peru 4.6K livinginperu.com
Philippines 25.1K sunstar.com.ph, journal.com.ph, bworldonline.com, newsbytes.ph, mindanews.com, tribwekchron.com, philstar.com
Qatar 8.8K aljazeera.com, middle-east-online.com
Romania 13.3K zmescience.com
Saint Kitts and Nevis 4.6K thestkittsnevisobserver.com
Saudi Arabia 42.8K arabnews.com, saudigazette.com.sa
Singapore 112.4K straitstimes.com
Somalia 197 mareeg.com
Somaliland 4.7K somalilandpress.com
South Africa 22.9K itweb.co.za, memeburn.com, themediaonline.co.za, news24.com, iafrica.com, mybroadband.co.za
South Korea 22K koreatimes.co.kr, yonhapnews.co.kr
Sri Lanka 33.8K lankabusinessonline.com, onlanka.com, lankanewspapers.com, groundviews.org
Tanzania 7.6K thecitizen.co.tz
Thailand 11.2K pattayamail.com
Trinidad 3.2K trinidadexpress.com
Turkey 2.5K theminaretonline.com, nationalturk.com, melodika.net
Uganda 6.7K monitor.co.ug
United Arab Emirates 108.8K emirates247.com, gulfnews.com, ameinfo.com, meed.com, 7days.ae
Venezuela 3.9K venezuelanalysis.com
Zambia 7.4K lusakatimes.com
Zimbabwe 26.1K newsday.co.zw, nehandaradio.com, thezimbabwemail.com
Table 27: News domains from underrepresented countries in RealNews.
I.1 Datasets
Aside from ToI Headlines, we use the original datasets, randomly subsampling as mentioned in Table 1.
I.2 Corpora